OpenStudy (anonymous):

simplify the square root of 16a^8b^-2

OpenStudy (saifoo.khan):

\[\Large \frac{\sqrt{16a^8}}{\sqrt {b^2}} \to \frac{4a^2}{b}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (saifoo.khan):

Welcome. :D